4000 A photon is used to break the iodine molecule, then the % of energy converted to the K.E. of iodine atoms if bond dissociation energy of I2 molecule is 246.5 kJ/mol


        (A)    8%


        (B)    12%


        (C)    17%


        (D)    25%


 


Grade:upto college level

2 Answers

Aditi Chauhan
10 years ago

(C)

 

Sol.   Energy of one photon =12400/4000

 

                                        = 3.1 ev

 

        Energy supplied by one ole photon in KJ/mole = 3.1 × 1.6 × 10–19 × 6 × 1023 × 10–3 = 297 kJmol–1

 

        % of energy converted to K.E. =  = 17%

Krishi Patel
2 years ago
Calculate the work function for metal for which ejection of the photoelectron can be stopped by applying 0.4  V  when the radiation of 300nm is used.